THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 110 — 2 April 2011

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Figure 1: The Tarantula Nebula (Pellegrini, Baldwin & Ferland).

Figure 2: Spectra of 29 newly confirmed quasars behind the SMC (Kozłowski, Kochanek & Udalski). 2 Figure 3: Could M 31 come from a major merger and eject the LMC away? (Fouquét et al. arXiv:1101.3748)

Editorial

Dear Colleagues, It is my pleasure to present you the 110^{th} issue of the Magellanic Clouds Newsletter. The Tarantula Nebula and star clusters, star formation, energetic phenomena and the broader galactic context – this issue is as diverse as ever. If you are looking for a job, check out the exciting opportunity at the University of Western Sydney! Hope you like the “cover” pictures – many thanks to the contributors. Pictures for next issue(s) are very welcome. The next issue is planned to be distributed in early June 2011.

Editorially Yours, Jacco van Loon

Refereed Journal Papers

Age and metallicity of star clusters in the Small Magellanic Cloud from integrated spectroscopy

Bruno Dias^{1}, Paula Coelho^{2}, Beatriz Barbuy^{1}, Leandro Kerber^{1,3} and Thais Idiart^{1}

1Universidade de São Paulo, Dept. de Astronomia, Rua do Matão 1226, São Paulo 05508-090, Brazil 2Núcleo de Astrofísica Teórica – Universidade Cruzeiro do Sul, R. Galvão Bueno, 868, sala 105, Liberdade, 01506-000, São Paulo, Brazil 3Universidade Estadual de Santa Cruz, Rodovia Ilhéus-Itabuna km16, 45662-000 Ilhéus, Bahia, Brazil

Context: Analysis of ages and metallicities of star clusters in the Magellanic Clouds provide information for studies on the chemical evolution of the Clouds and other dwarf irregular galaxies. Aims: The aim is to derive ages and metallicities from integrated spectra of 14 star clusters in the Small Magellanic Cloud, including a few intermediate/old age star clusters. Methods: Making use of a full-spectrum fitting technique, we compared the integrated spectra of the sample clusters to three different sets of single stellar population models, using two fitting codes available in the literature. Results: We derive the ages and metallicities of 9 intermediate/old age clusters, some of them previously unstudied, and 5 young clusters. Conclusions: We point out the interest of the newly identified as intermediate/old age clusters HW 1, NGC 152, Lindsay 3, Lindsay 11, and Lindsay 113. We also confirm the old ages of NGC 361, NGC 419, Kron 3, and of the very well-known oldest SMC cluster, NGC 121.

Published in Astronomy & Astrophysics, 520A, 85 (2010) Available from arXiv:1002.4301

Structure and feedback in 30 Doradus II. Structure and chemical abundances

E.W. Pellegrini^{1}, J.A. Baldwin^{2} and G.J. Ferland^{3}

1Department of Astronomy, University of Michigan, 500 Chruch Street, Ann Arbor, MI, 48109, USA 2Physics and Astronomy Department, Michigan State University, 3270 Biomedical Physical Sciences Building, East Lansing, MI, 48824, USA 3Department of Physics and Astronomy, University of Kentucky, 177 Chemistry/Physics Building, Lexington, KY 40506, USA

We use our new optical-imaging and spectrophotometric survey of key diagnostic emission lines in 30 Doradus, together with cloudy photoionization models, to study the physical conditions and ionization mechanisms along over 4000 individual lines of sight at points spread across the face of the extended nebula, out to a projected radius 75 pc from R 136 at the center of the ionizing cluster NGC 2070. We focus on the physical conditions, geometry and importance of radiation pressure on a point-by-point basis, with the aim of setting observational constraints on important feedback processes. The completeness of our survey also allows us to create a composite spectrum of 30 Doradus, simulating the observable spectrum of a spatially-unresolved, distant giant extragalactic H ii region. With these data we test various strong line abundance techniques that are typically applied to an entire giant extragalactic H ii regions or galaxies, by comparing them to the abundance measurements made from our spatially resolved observations. We find that the abundances and physical parameters derived from the globally averaged optical spectrum using the strong-line methods agree with our abundances measured from the spatially resolved data. This is in spite of the more than one order of magnitude range in the ionization parameter and density of the actual gas in 30 Dor.

Submitted to Astrophysical Journal Available from arXiv:1101.3778 and from http://astro.lsa.umich.edu/∼pelleger/

XMM-Newton observations of the superbubble in N 158 in the LMC

Manami Sasaki^{1}, Dieter Breitschwerdt^{2}, Verena Baumgartner^{3} and Frank Haberl^{4}

1Institute for Astronomy and Astrophysics, University of Tübingen, Germany 2Department of Astronomy and Astrophysics, Berlin Institute of Technology, Germany 3Institute for Astronomy, University of Vienna, Austria 4Max-Planck-Institute for Extraterrestrial Physics, Garching, Germany

Aims: We study the diffuse X-ray emission observed in the field of view of the pulsar B 0540−69 in the Large Magellanic Cloud (LMC) by XMM-Newton. We want to understand the nature of this soft diffuse emission, which coincides with the superbubble in the H ii region N 158, and improve our understanding of the evolution of superbubbles. Methods: We analyse the XMM-Newton spectra of the diffuse emission. Using the parameters obtained from the spectral fit, we perform calculations of the evolution of the superbubble. The mass loss and energy input rates are based on the initial mass function (IMF) of the observed OB association inside the superbubble. Results: The analysis of the spectra shows that the soft X-ray emission arises from hot shocked gas surrounded by a thin shell of cooler, ionised gas. We show that the stellar winds alone cannot account for the energy inside the superbubble, but the energy release of 2–3 supernova explosions in the past ∼ 1 Myr provides a possible explanation. Conclusions: The combination of high sensitivity X-ray data, allowing spectral analysis, and analytical models for superbubbles bears the potential to reveal the evolutionary state of interstellar bubbles, if the stellar content is known.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:1010.0783

An X-ray investigation of three supernova remnants in the Large Magellanic Cloud

Matthew D. Klimek^{1}, S.D. Points^{1}, R.C. Smith^{1}, R.L. Shelton^{2} and R. Williams^{3}

1Cerro Tololo Inter-American Observatory, National Optical Astronomy Observatory, Casilla 603, La Serena, Chile 2Department of Physics and Astronomy, University of Georgia, Athens, GA 30602, USA 3Columbus State University, Coca-Cola Space Science Center, 701 Front Ave., Columbus, GA 31901, USA

We have investigated three SNRs in the LMC using multi-wavelength data. These SNRs are generally fainter than the known sample and may represent a previously missed population. One of our SNRs is the second LMC remnant analyzed which is larger than any Galactic remnant for which a definite size has been established. The analysis of such a large remnant contributes to the understanding of the population of highly evolved SNRs. We have obtained X-ray images and spectra of three of these recently identified SNRs using the XMM-Newton observatory. These data, in conjunction with pre-existing optical emission-line images and spectra, were used to determine the physical conditions of the optical- and X-ray-emitting gas in the SNRs. We have compared the morphologies of the SNRs in the different wavebands. The physical properties of the warm ionized shell were determined from the Hα surface brightness and the SNR expansion velocity. The X-ray spectra were fit with a thermal plasma model and the physical conditions of the hot gas were derived from the model fits. Finally, we have compared our observations with simulations of SNR evolution.

Published in ApJ, 725, 2281 (2010) Available from arXiv:1101.5669

An interesting candidate for isolated massive star formation in the Small Magellanic Cloud

R. Selier^{1}, M. Heydari-Malayeri^{1} and D.A. Gouliermis^{2}

1LERMA, Observatoire de Paris, France 2Max-Planck-Institut für Astronomie, Heidelberg, Germany

The SMC region with which this paper is concerned has the particularity of containing the highest concentration of IRAS/Spitzer sources, H i emission, and molecular clouds in this neighboring galaxy. However very few studies have been devoted to it, despite these signs of star formation. We present the first detailed study of the compact H ii region N 33 in the SMC by placing it in a wider context of massive star formation. Moreover, we show that N 33 is a particularly interesting candidate for isolated massive star formation. This analysis is based mainly on optical ESO NTT observations, both imaging and spectroscopy, coupled with other archive data, notably Spitzer images (IRAC 3.6, 4.5, 5.8, and 8.0 μm) and 2MASS observations. We derive a number of physical characteristics of the compact H ii region N 33 for the first time. This gas and dust formation of 7.4′′ (2.2 pc) in diameter is powered by a massive star of spectral type O6.5–O7 V. The compact H ii region belongs to a rare class of H ii regions in the Magellanic Clouds, called high-excitation blobs (HEBs). We show that this H ii region is not related to any star cluster. Specifically, we do not find any traces of clustering around N 33 at scales larger than 10′′ (∼ 3 pc). On smaller scales there is a marginal stellar concentration, the low density of which, below the 3-σ level, does not classify it as a real cluster. We also verify that N 33 is not a member of any large stellar association. Under these circumstances, N 33 is therefore attractive also because it represents a remarkable case of isolated massive star formation in the SMC. Various aspects of the relevance of N 33 to the topic of massive star formation in isolation are discussed.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1102.1257

On interpretation of recent proper motion data for the Large Magellanic Cloud

Kenji Bekki^{1}

1ICRAR at UWA, Australia

Recent observational studies using the Hubble Space Telescope (HST) have derived the center-of-mass proper motion (CMPM) of the Large Magellanic Cloud (LMC). Although these studies carefully treated both rotation and perspective effects in deriving the proper motion for each of the sampled fields, they did not consider the effects of local random motion in the derivation. This means that the average PM of the fields (i.e., the observed CMPM) could significantly deviate from the true CMPM, because the effect of local random motion can not be close to zero in making the average PM for the small number of the fields (∼ 10). We discuss how significantly the observationally derived CMPM can deviate from the true CMPM by applying the same method as used in the observations for a dynamical model of the LMC with a known true CMPM. We find that the deviation can be as large as ∼ 50 km s^{−1} (∼ 0.21 mas yr^{−1}), if the LMC has a thick disk and a maximum circular velocity of ∼ 120 km s^{−1}. We also find that the deviation depends both on the total number of the sampled fields and on structure and kinematics of the LMC. We therefore suggest that there is a possibility that the observed CMPM of the LMC deviates from the true one to some extent. We also show that a simple mean of PM for a large number of the LMC fields (∼ 1000) can be much closer to the true CMPM.

Accepted for publication in ApJL Available from arXiv:1102.1504

The star formation history of the Large Magellanic Cloud star cluster NGC 1751

Stefano Rubele^{1,2}, Léo Girardi^{1}, Vera Kozhurina-Platais^{3}, Paul Goudfrooij^{3} and Leandro Kerber^{4}

1OA Padova, Italy 2Dip. Astro. Padova, Italy 3STScI Baltimore, USA 4UESC Ilhéus, Brazil

The HST/ACS colour–magnitude diagrams (CMD) of the populous LMC star cluster NGC 1751 present both a broad main sequence turn-off and a dual clump of red giants. We show that the latter feature is real and associate it to the first appearance of electron-degeneracy in the H-exhausted cores of the cluster stars. We then apply to the NGC 1751 data the classical method of star formation history (SFH) recovery via CMD reconstruction, for different radii corresponding to the cluster centre, the cluster outskirts, and the underlying LMC field. The mean SFH derived from the LMC field is taken into account during the stage of SFH-recovery in the cluster regions, in a novel approach which is shown to significantly improve the quality of the SFH results. For the cluster centre, we find a best-fitting solution corresponding to prolonged star formation for a for a timespan of 460 Myr, instead of the two peaks separated by 200 Myr favoured by a previous work based on isochrone fitting. Remarkably, our global best-fitting solution provides an excellent fit to the data – with χ^{2} and residuals close to the theoretical minimum – reproducing all the CMD features including the dual red clump. The results for a larger ring region around the centre indicate even longer star formation, but in this case the results are of lower quality, probably because of the differential extinction detected in the area. Therefore, the presence of age gradients in NGC 1751 could not be probed. Together with our previous findings for the SMC cluster NGC 419, the present results for the NGC 1751 centre argue in favour of multiple star formation episodes (or continued star formation) being at the origin of the multiple main sequence turn-offs in Magellanic Cloud clusters with ages around 1.5 Gyr.

Accepted for publication in MNRAS Available from arXiv:1102.2814

The extended main-sequence turn-off clusters of the Large Magellanic Cloud — Missing links in globular cluster evolution

Stefan C. Keller^{1}, A. Dougal Mackey^{1} and Gary S. Da Costa^{1}

1Research School of Astronomy and Astrophysics, Australian National University, Australia

Recent observations of intermediate age (1–3 Gyr) massive star clusters in the Large Magellanic Cloud (LMC) have revealed that the majority possess bifurcated or extended main-sequence turn-off (EMSTO) morphologies. This effect can be understood to arise from subsequent star formation amongst the stellar population with age differences between constituent stars amounting to 50–300 Myr. Age spreads of this order are similarly invoked to explain the light element abundance variations witnessed in ancient globular clusters. In this paper we explore the proposition that the clusters exhibiting the EMSTO phenomenon are a general phase in the evolution of massive clusters, one that naturally leads to the particular chemical properties of the ancient globular cluster population. We show that the isolation of EMSTO clusters to intermediate ages is the consequence of observational selection effects. In our proposed scenario, the EMSTO phenomenon is identical to that which establishes the light element abundance variations that are ubiquitous in the ancient globular cluster population. Our scenario makes a strong prediction: EMSTO clusters will exhibit abundance variations in the light elements characteristic of the ancient GC population.

Accepted for publication in ApJ Available from arXiv:1102.1723

The Magellanic Quasars Survey. I. Doubling the number of known AGNs behind the Small Magellanic Cloud

Szymon Kozłowski^{1,2}, Christopher S. Kochanek^{2,3} and Andrzej Udalski^{1}

1Warsaw University Observatory, Poland 2The Ohio State University, USA 3Center for Cosmology and Astroparticle Physics, USA

We report the spectroscopic confirmation of 29 new, 12 plausible, and 3 previously known quasars behind the central ∼ 1.5 deg^{2} region of the Small Magellanic Cloud. These were identified in a single 2df/AAΩ observation on the Anglo–Australian Telescope of 268 candidates selected primarily based on their mid-IR colors, along with a smaller number of optically variable sources in OGLE-II close to known X-ray sources. The low detection efficiency was partly expected from the high surface density of SMC as compared to the LMC targets and the faintness of many of them (149 with I > 20 mag). The expected number of I < 20 mag quasars in the field is ∼ 38, and we found 15 (22 with plausible) or 40% (60%). We did not attempt to determine the nature of the remaining sources, although several appear to be new planetary nebulae. The newly discovered AGNs can be used as reference points for future proper motion studies, to study absorption in the interstellar medium, and to study the physics of quasar variability with the existing long-term, highly cadenced OGLE light curves.

Submitted to ApJ Available from arXiv:1102.0703

The Swift-BAT survey reveals the orbital period of three high-mass X-ray binaries

A. D’Aì^{1}

1Dipartimento di Fisica, University of Palermo, Italy

A growing number of previously hidden Galactic X-ray sources are now detected with recent surveys performed by the Integral and Swift satellites. Most of these new sources eluded past surveys due to their large local X-ray extinction and consequent low soft X-ray flux. The Swift-BAT performs daily monitoring of the sky in an energy band (15–150 keV) which is only marginally affected by X-ray extinction, thus allowing for the search of long periodicities in the light curve and identification of the nature of the X-ray sources. We performed a period search using the folding technique in the Swift-BAT light curves of three Integral sources: IGR J05007−7047, IGR J13186−6257 and IGR J17354−3255. Their periodograms show significant peaks at 30.77 ± 0.01 d, 19.994 ± 0.01 d and 8.448 ± 0.002 d, respectively. We estimate the significance of these features from the χ^{2} distribution of all the trials, finding a probability < 1.5 × 10^{−4} that the detections occurred due to chance. We complement our analysis with the study of their broadband X-ray emission. We identify the periodicities with the orbital periods of the sources. The periods are typical for the wind accretors X-ray binaries and we support this identification showing that also their energy spectra are compatible with an X-ray spectral emission characteristic of high-mass X-ray binaries. The spectrum of IGR J05007−704 that resides in the Large Magellanic Cloud, does not show any intrinsic local absorption, whereas the spectra of the Galactic sources IGR J17354−3255 and IGR J13186−6257 may be affected by a local absorber. The folded light curve for IGR J13186−6257 suggests a possible Be companion star.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1102.4546

Improved prospects for the detection of new Large Magellanic Cloud planetary nebulae

Brent Miszalski^{1}, Ralf Napiwotzki^{1}, Maria-Rosa L. Cioni^{1,2} and Jundan Nie^{3,4}

1Centre for Astrophysics Research, STRI, University of Hertfordshire, College Lane Campus, Hatfield AL10 9AB, UK 2University Observatory Munich, Scheinerstraße 1, D-81679, München, Germany 3Research School of Astronomy and Astrophysics, Australian National University, Cotter Road, Weston Creek ACT 2611, Australia 4Department of Astronomy, Beijing Normal University, Beijing, 100875, China

The Large Magellanic Cloud (LMC) contains the nearest large extragalactic population of planetary nebulae (PNe). A shallow viewing angle and low interstellar reddening towards the LMC potentially means a larger, more complete flux-limited population can be assembled than for any other galaxy. These advantages appear to be reflected by the small gap between the catalogued (∼ 700 PNe) and estimated (1000 ± 250 PNe) population size. With more detailed multi-wavelength studies the catalogued number of LMC PNe may fall, potentially widening this gap. We demonstrate here that the gap can be further bridged with improved optical and near-infrared imaging surveys. We present three [O iii]-selected PNe discovered from ESO WFI observations of the 30 Doradus region and one serendipitous discovery from near-infrared VISTA Magellanic Cloud (VMC) survey observations. The WFI PNe have resolved [O iii] and Hα nebulae that verify their PN nature and their [O iii] fluxes place them 6–7 mag (m_{5007} = 20–21 mag) fainter than the bright-end of the planetary nebula luminosity function (PNLF). Their faintness, small angular size and surrounding complex emission-line background explains why previous Hα surveys of the region did not select them. We estimate there may be as many as 50–75 similar PNe awaiting discovery in the central 5^{∘} × 5^{∘} of the LMC. The VMC survey routinely detects PNe as red resolved nebulae that may allow some of this expected population to be recovered without traditional narrow-band imaging surveys. We demonstrate this potential with the first new VMC-selected PN which has a rare Wolf–Rayet [WC9]–[WC11] central star.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:1103.1957

Small-scale structure in the SDSS and ΛCDM: Isolated L_{⋆} galaxies with bright satellites

Erik J. Tollerud^{1}, Michael Boylan-Kolchin^{1,2}, Elizabeth J. Barton^{1}, James S. Bullock^{1} and Christopher Q. Trinh^{3,1}

1University of California, Irvine, USA 2Max-Planck-Institut für Astrophysik, Garching bei München, Germany 3University of Sydney, Australia

We use a volume-limited spectroscopic sample of isolated galaxies in the Sloan Digital Sky Survey (SDSS) to investigate the frequency and radial distribution of luminous (M_{r}< ∼−18.3) satellites like the Large Magellanic Cloud (LMC) around ∼L_{⋆} Milky Way analogs and compare our results object-by-object to ΛCDM predictions based on abundance matching in simulations. We show that 12% of Milky Way-like galaxies host an LMC-like satellite within 75 kpc (projected), and 42% within 250 kpc (projected). This implies ∼ 10% have a satellite within the distance of the LMC, and ∼ 40% of L_{⋆} galaxies host a bright satellite within the virialized extent of their dark matter halos. Remarkably, the simulation reproduces the observed frequency, radial dependence, velocity distribution, and luminosity function of observed secondaries exceptionally well, suggesting that ΛCDM provides an accurate reproduction of the observed Universe to galaxies as faint as L ∼ 10^{9} L_{⊙} on ∼ 50 kpc scales. When stacked, the observed projected pairwise velocity dispersion of these satellites is σ ∼ 160 km s^{−1}, in agreement with abundance-matching expectations for their host halo masses. Finally, bright satellites around L_{⋆} primaries are significantly redder than typical galaxies in their luminosity range, indicating that environmental quenching is operating within galaxy-size dark matter halos that typically contain only a single bright satellite. This redness trend is in stark contrast to the Milky Way’s LMC, which is unusually blue even for a field galaxy. We suggest that the LMC’s discrepant color might be further evidence that it is undergoing a triggered star-formation event upon first infall.

Submitted to Astrophysical Journal Available from arXiv:1103.1875

The XMM-Newton survey of the Small Magellanic Cloud: XMMU J005011.2−730026 = SXP 214, a Be/X-ray binary pulsar

M.J. Coe on behalf of the XMM SMC Survey Team^{1}

1University of Southampton, UK

In the course of the XMM-Newton survey of the Small Magellanic Cloud (SMC), a region to the east of the emission nebula N 19 was observed in November 2009. To search for new candidates for high mass X-ray binaries the EPIC PN and MOS data of the detected point sources were investigated and their spectral and temporal characteristics identified. A new transient (XMMU J005011.2−730026 = SXP 214) with a pulse period of 214.05 s was discovered; the source had a hard X-ray spectrum with power-law index of ∼ 0.65. The accurate X-ray source location permits the identification of the X-ray source with a ∼ 15^{th} magnitude Be star, thereby confirming this system as a new Be/X-ray binary.

Accepted for publication in MNRAS Available from arXiv:1103.0195

The VLT-FLAMES Tarantula Survey I: Introduction and observational overview

C.J. Evans & 41 co-authors^{1}

1UK ATC, UK

The VLT-FLAMES Tarantula Survey (VFTS) is an ESO Large Programme that has obtained multi-epoch optical spectroscopy of over 800 massive stars in the 30 Doradus region of the Large Magellanic Cloud (LMC). Here we introduce our scientific motivations and give an overview of the survey targets, including optical and near-infrared photometry and comprehensive details of the data reduction. One of the principal objectives was to detect massive binary systems via variations in their radial velocities, thus shaping the multi-epoch observing strategy. Spectral classifications are given for the massive emission-line stars observed by the survey, including the discovery of a new Wolf–Rayet star (VFTS 682, classified as WN5h), 2′ to the northeast of R 136. To illustrate the diversity of objects encompassed by the survey, we investigate the spectral properties of sixteen targets identified by Gruendl & Chu from Spitzer photometry as candidate young stellar objects or stars with notable mid-infrared excesses. Detailed spectral classification and quantitative analysis of the O- and B-type stars in the VFTS sample, paying particular attention to the effects of rotational mixing and binarity, will be presented in a series of future articles to address fundamental questions in both stellar and cluster evolution.

Accepted for publication in A&A Available from arXiv:1103.5386

The VLT-FLAMES Tarantula Survey II: R 139 revealed as a massive binary system

W.D. Taylor^{1}, C.J. Evans, H. Sana, N.R. Walborn, S.E. de Mink, V.E. Stroud, A. Alvarez-Candal, R.H. Barba,

J.M. Bestenlehner, A.Z. Bonanos, I. Brott, P.A. Crowther, A. de Koter, K. Friedrich, G. Gräfener, V. Hénault-Brunet, A. Herrero, L. Kaper, N. Langer, D.J. Lennon, J. Maíz Appellániz, N. Markova, N. Morrell, L.

Monaco and J.S. Vink

1Institute for Astronomy, Royal Observatory Edinburgh, UK

We report the discovery that R 139 in 30 Doradus is a massive spectroscopic binary system. Multi-epoch optical spectroscopy of R 139 was obtained as part of the VLT-FLAMES Tarantula Survey, revealing a double-lined system.

The two components are of similar spectral types; the primary exhibits strong C iii 4650 emission and is classified as an O6.5 Iafc supergiant, while the secondary is an O6 Iaf supergiant. The radial-velocity variations indicate a highly eccentric orbit with a period of 153.9 days. Photometry obtained with the Faulkes Telescope South shows no evidence for significant variability within an 18 month period. The orbital solution yields lower mass limits for the components of M_{1} sin^{3} i = 78 ± 8 M_{⊙} and M_{2} sin^{3} i = 66 ± 7 M_{⊙}. As R 139 appears to be the most massive binary system known to contain two evolved Of supergiants, it will provide an excellent test for atmospheric and evolutionary models.

Accepted for publication in A&A Available from arXiv:1103.5387

Conference Paper

Iron depletion in ionized nebulae of the Large Magellanic Cloud

G. Delgado-Inglada^{1}, M. Rodríguez^{1}, J. García-Rojas^{2}, M. Peña^{3} and M.T. Ruiz^{4}

1Instituto Nacional de Astrofísica, Optica y Electrónica (INAOE), Apdo. Postal 51 y 216, 72000 Puebla, México 2Instituto de Astrofísica de Canarias (IAC), C/ Vía Láctea s/n, E38205 La Laguna, Tenerife, Spain 3Instituto de Astronomía, Universidad Nacional Autónoma de México, México 4Departamento de Astronomía, Universidad de Chile, Chile

We present here preliminary results of an analysis of the iron abundance in the ionized gas of three planetary nebulae and one H ii region of the Large Magellanic Cloud (LMC). These results are compared with the ones we obtain for a sample of Galactic and extragalactic nebulae. We find that the amounts of iron depletion into dust grains in LMC nebulae are similar to those found in Galactic nebulae. Objects with lower metallicities show lower depletions, but a larger sample of objects is needed to explore the reasons behind this trend.

Oral contribution, published in ”XII Latin American Regional IAU Meeting”, Rev. Méx. Astron. and Astrofis. Conf. Ser. Available from arXiv:1103.2158

Job Advert

Lecturer/Senior Lecturer in Computational Imaging Visual Science and Computational Astrophysics School of Computing and Mathematics University of Western Sydney, Australia

3 Years Fixed Term initially with opportunity to be converted to Permanent.

Applications are invited for the position of a Lecturer (Level B) or Senior Lecturer (Level C) in Computational Imaging, Visual Science and Computational Astrophysics to complement the existing research expertise in the School. The successful applicant will be an enthusiastic and experienced scholar who will be expected to take a major role in the School in leading the development of teaching and research programs that complement and expand the School’s existing programs in the areas of Computer Science, Information and Communications Technology (ICT), Applied Mathematics and Computational Astrophysics.

We seek to appoint an energetic, forward thinking, dynamic, creative and innovative interdisciplinary scholar in Computational Imaging, Visual Science and Computational Astrophysics, who brings together scholarship that covers science imaging, broad computational astrophysics and image analysis, as well as experience in data and information visualisation, visual computing and visual data mining.

Applicants must hold a PhD in Computing Science, Information Technology or a closely related field, such as Computational Astrophysics, Scientific Computing, Computational Science, Image Analysis, Design Computing, Design and Computation, or a closely related field, and be able to demonstrate the capacity to play a significant role in scholarship, research, and professional development activities. This role will involve active engagement in the extension of existing and the development of a new program of research and supervision of higher degree research students. The successful applicant will have a background and experience appropriate to the leadership role envisaged.

Applicants must address the selection criteria specified in the Position Statement, which is available from the University Website, pointing out the specific level – B or C, for which they apply. For information about the School please visit the UWS website: http://www.uws.edu.au/computing mathematics/scm and the School website: http://www.scm.uws.edu.au/.

Remuneration Package from 13 May 2011: Lecturer (Academic Level B) $96,915 to $114,361 p.a. (comprising Salary of $81,894 to $96,703 p.a.; 17% Superannuation, and Leave Loading); Senior Lecturer (Academic Level C) $117,823 to $135,155 p.a. (comprising Salary of $99,662 to $114,476 p.a.; 17% Superannuation, and Leave Loading).

Position Enquiries: Professor Simeon Simoff, +61 (2) 9685 9179, s.simoff@uws.edu.au

Closing Date: 29 April 2011

See also https://uws.nga.net.au/?jati=772A1B8B-C215-9177-C6B0-6457FF9A110B